OpenStudy (anonymous):

Identify the y-intercept. y = -1.5x – 33

OpenStudy (anonymous):

-33

OpenStudy (anonymous):

y= -1.5 (0) -33 =-33

OpenStudy (anonymous):

-33

OpenStudy (anonymous):

Compare it with y=mx +c where c =intercept so c= -33